mathslover (mathslover):

Differentiation Problem (4)... Find \(\cfrac{dy}{dx}\) for the equation\(x^2y^3 + y\sin x = 0\)

mathslover (mathslover):

I know that we have to use implicit differentiation here.

mathslover (mathslover):

And, yes, I just studied it.

Miracrown (miracrown):

Good start !

mathslover (mathslover):

So, before anyone wastes his/her time, I will write that out.

mathslover (mathslover):

On differentiating that, I get : \(2x y^3 + x^2 \left( 3y^2 \cfrac{dy}{dx} \right) + y \cos x + \cfrac{dy}{dx} \sin x = 0 \)

mathslover (mathslover):

Or just using the shortcut : \(\cfrac{dy}{dx} = \cfrac{-(2xy^3 + y\cos x) }{3x^2 y^2 + \sin x}\) The shortcut I used : \(\cfrac{dy}{dx} = \cfrac{\text{Derivative of function w.r.t x (taking y as constant)}}{\text{Derivative of function w.r.t y( taking x as constant) }}\)

mathslover (mathslover):

Am I correct till now?

Miracrown (miracrown):

Yep, you're correct!

mathslover (mathslover):

Okay, I think I am getting something now : From the question : \(x^2 y^3 + y\sin x = 0\) \(x^2 y^3 = -y \sin x\) \(x^2 y^2 = - \sin x\)

mathslover (mathslover):

\(\cfrac{dy}{dx} = \cfrac{-(2xy^3 + y\cos x) }{3x^2 y^2 + \sin x}\) becomes : \(\cfrac{dy}{dx} = \cfrac{-(2xy^3 + y\cos x) }{-3\sin x + \sin x} \\ \cfrac{dy}{dx} = \cfrac{-(2xy^3 + y\cos x) }{-2\sin x} \)

mathslover (mathslover):

Or : \(\cfrac{dy}{dx} = \cfrac{\cancel{-}(2xy^3 + y\cos x) }{\cancel{-}2\sin x}\)

mathslover (mathslover):

\(\cfrac{dy}{dx} = \cfrac{2xy^3}{2\sin x} + \cfrac{y\cos x}{ 2\sin x}\) \(\cfrac{dy}{dx} = {xy^3 \times \csc x} + \cfrac{y}{2} \cot x\)

mathslover (mathslover):

That seems to be the finally answer, right? Or is there any simplification required?

Miracrown (miracrown):

Why are you changing x^2 y^2 to -sin x?

mathslover (mathslover):

Because it is given in the question..

Miracrown (miracrown):

This first one is the final answer

mathslover (mathslover):

Did you mean to say that the one without that simplificatin is the final answer?

mathslover (mathslover):

*simplification..

mathslover (mathslover):

\(\color{blue}{\text{Originally Posted by}}\) @mathslover \(\cfrac{dy}{dx} = \cfrac{-(2xy^3 + y\cos x) }{3x^2 y^2 + \sin x}\) \(\color{blue}{\text{End of Quote}}\) This one?

Miracrown (miracrown):

Yes, I think so. I don't think your simplification is valid

Miracrown (miracrown):

Since 3x^2 y^2 is not equal to -3sin x

Miracrown (miracrown):

Oh, never mind! I'm so sorry!

Miracrown (miracrown):

I see what you were doing now. My bad! >.<

mathslover (mathslover):

Well, yeah, the condition stands TALL : \(y \ne 0\) ... if y = 0 I can not say that x^2y^2 = -sin x

Miracrown (miracrown):

see this is why not getting enough sleep does to you

Miracrown (miracrown):

Yep, that's all correct!

Miracrown (miracrown):

Although the answer will still work You can check it manually And if y = 0, then dy/dx = 0

mathslover (mathslover):

Well yeah... why don't u have 2-3 days of continuou sleep... may be, this will help in future :P

mathslover (mathslover):

But @Miracrown -> if y = 0 then I can NOT cancel out y in the equation x^2 y^3 + y sin x = 0 so, basically : y should not be zero for the final condition I got - \(\color{blue}{\text{Originally Posted by}}\) @mathslover \(\cfrac{dy}{dx} = \cfrac{2xy^3}{2\sin x} + \cfrac{y\cos x}{ 2\sin x}\) \(\cfrac{dy}{dx} = {xy^3 \times \csc x} + \cfrac{y}{2} \cot x\) \(\color{blue}{\text{End of Quote}}\)

Miracrown (miracrown):

Unless x=0 too, and then who knows?

Miracrown (miracrown):

Yeah, you're right but there's a workaround Take the very first equation you have -- the implicit one And set y=0 in there

mathslover (mathslover):

Yeah...! So, if my final answer is \(\color{blue}{\text{Originally Posted by}}\) @mathslover \(\cfrac{dy}{dx} = \cfrac{2xy^3}{2\sin x} + \cfrac{y\cos x}{ 2\sin x}\) \(\cfrac{dy}{dx} = {xy^3 \times \csc x} + \cfrac{y}{2} \cot x\) ------ FINAL \(\color{blue}{\text{End of Quote}}\) Then , I will have to say that y should not be equal to zero. Right?

Miracrown (miracrown):

You get that dy/dx sin x = 0 So dy/dx = 0 unless sin x = 0 And in that case, it's undefined

Miracrown (miracrown):

So it works for that case even though the work didn't You can say that EITHER y can't be 0 OR sin x can't be 0

mathslover (mathslover):

dy/dx = 0 , then doesn't it mean that y = constant? So , why is it undefined?

OpenStudy (kainui):

Out of curiosity, what if you didn't do implicit differentiation, divided out one of the y and solved the quadratic of y, then took the derivative? Does that mean there are now two derivatives and we're missing one some how?

Miracrown (miracrown):

y is only "constant" in very small areas

Miracrown (miracrown):

Well, we should get the same answer for the derivative If we did it correctly

Miracrown (miracrown):

divide by y, and we get: x^2 y^2 + sin x = 0 (Like before, we can only do this if y isn't 0) Then x^2 y^2 = -sin x So y^2 = -sin x/x^2 y = +/- sqrt(-sin x/x^2)

mathslover (mathslover):

Clearly : \(y^2 = \cfrac{-\sin x}{x^2}\)

mathslover (mathslover):

Yeah... mira got it..

Miracrown (miracrown):

Then we take the derivative of this dy/dx = +/- 1/ sqrt(-sin x/x^2) * (-cos x x^2 + 2x sin x)/x^4

mathslover (mathslover):

We have "-" inside the square roots, wouldn't that turn into something imaginary?

mathslover (mathslover):

Oh.. sorry, got it.

Miracrown (miracrown):

And apparently that's the same thing

Miracrown (miracrown):

It would -- if sin x < 0 Good catch But the same thing happens in the original equation Look at the original equation -- if sin x < 0

Miracrown (miracrown):

Sorry -- all of this is a problem if sin x > 0 Then x^2 y^3 will have the same sign as y sin x So their sum can't be 0 Which means it's undefined if sin x > 0 Basically, this is a weird problem

OpenStudy (kainui):

Oooh pretty. http://www.wolframalpha.com/input/?i=y%3Dsqrt%28-sinx%29%2Fx

Miracrown (miracrown):

Yeah, that looks cool!

Miracrown (miracrown):

But crazy too You can see it's imaginary a lot of the time

mathslover (mathslover):

\[\cfrac{dy}{dx} = \pm \cfrac{1}{\sqrt{ \left( \cfrac{-sin x}{x^2} \right) }} \times \cfrac{(-cos x \times x^2 + 2x sin x)}{x^4} \]

mathslover (mathslover):

Is this what you meant @Miracrown ...

mathslover (mathslover):

I must appreciate you all... You all have done well in confusing me :P

OpenStudy (kainui):

Actually if you graph it it'll look like this: